# On the Non-Identifiability of Steering Vectors in Large Language Models

### 05 Mar 2026 v3

@@ 7 Limitations
Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor). While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions. Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human evaluation, providing a conservative test of observational equivalence.
Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization or invariance objectives (IRM) can recover identifiable representations. Empirical validation of these approaches remains an important direction for future work. Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
8 Future Work
Our results establish the necessity of validation beyond input-output behavior. Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components and characterizing how null-space dimensionality scales with model size.
Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability. Human evaluation would provide stronger validation than our lexical heuristics and direct measurement of null-space dimensionality would quantify our theoretical predictions.
@@ 9 Conclusion
We have established that steering vectors are fundamentally non-identifiable under the local linear approximation, with infinitely many distinct directions producing equivalent behavioral effects. This non-identifiability persists across models, traits and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may fail to yield interpretable representations. Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone. More broadly, when internal interventions admit large equivalence classes under the model’s local geometry, consistent output shifts do not imply that a discovered direction corresponds to a canonical or causally privileged feature. Stronger interpretability claims therefore require structural constraints or additional assumptions that break these symmetries and distinguish meaningful factors from geometrically arbitrary ones.
@@ Acknowledgments
We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments. We also acknowledge the HuggingFace platform and maintainers of the open-source models used in this work. Generative AI tools were used for language editing and code assistance; all technical content was reviewed and verified by the authors.
Appendix A Logit-Level Equivalence Test
To complement our semantic evaluations, we perform a direct logit-level equivalence test. For each model and trait, we compare the next-token logit vectors produced by steering with $v$ and with $v+v_{\perp}$, where $v_{\perp}$ is a norm-matched direction orthogonal to $v$. We measure the $\ell_{2}$ distance between logit vectors and normalize by the corresponding distance induced by a norm-matched random direction:
@@ Appendix A Logit-Level Equivalence Test
Values $R<1$ indicate that orthogonal perturbations induce smaller logit deviations than random directions. In all tested models and traits, this condition holds. Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions. The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$ and top-10 token overlap ranges from $86\%$ to $96\%$.
These results provide direct evidence that equivalence is not an artifact of semantic measurement. While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes. Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement and top-$k$ overlap for all models and traits.
Table 3: Logit-level equivalence statistics. $R$ denotes the ratio $v_{\perp}$ / random. Agreement% refers to token agreement, and Overlap% indicates top-10 overlap.
Model Trait $R$ Agreement% Overlap%
Qwen-3B Humor 0.90 81% 86%
Formality 0.76 94% 96%
Politeness 0.75 93% 96%
Llama-8B Humor 0.64 96% 90%
Formality 0.49 93% 95%
Politeness 0.65 91% 94%
Appendix B Detailed Proof of Proposition 1
B.1 Null-Space Ambiguity (Constructive Proof)
Setup Consider the local linear approximation of the steering effect:
@@ B.1 Null-Space Ambiguity (Constructive Proof)
where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian.
Step 1: Null space characterization
@@
Step 2: Rank Bound
The Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ has maximum possible rank:
$\text{rank}(J_{\ell})\leq\min(V,d)$ (14)
In modern language models, the hidden dimension is typically $d\approx 4000$ while the vocabulary size is $V\approx 50000$. Therefore, the Jacobian satisfies $\max\mathrm{rank}(J_{\ell})=d$.
Step 3: Effective rank is much lower
Output distributions lie on a low-dimensional manifold. The effective rank satisfies:
$\text{rank}_{\epsilon}(J_{\ell})=\#\{\sigma_{i}:\sigma_{i}>\epsilon\cdot\sigma_{\max}\}\ll d$ (15)
where $\sigma_{i}$ are singular values of $J_{\ell}$ and $\epsilon$ is a threshold (e.g., $10^{-4}$).
Intuition: Empirical evidence suggests that in overparameterized LLMs the effective rank of $J_{\ell}$ may be substantially lower than $d$, consistent with low dimensional structure observed in related weight space studies (Li et al., 2018; Maennel et al., 2018) and concentration of output sensitivity in a limited number of directions (Ghorbani et al., 2019). Consequently, $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is often positive in practice. However, this empirical observation is not required for the proof; the argument establishes non identifiability whenever $\dim(\mathcal{N})\geq 1$.
Step 4: Constructing equivalent vectors
For any steering vector $v\in\mathbb{R}^{d}$ and any $v_{0}\in\mathcal{N}$, define
$v^{\prime}=v+v_{0}.$ (16)
Then for all $x$ and all $\alpha$:
$J_{\ell}(x)v^{\prime}$ $=J_{\ell}(x)(v+v_{0})$ (17)
$=J_{\ell}(x)v+J_{\ell}(x)v_{0}$
$=J_{\ell}(x)v.$
Therefore:
$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (18)
$=o(x,0,0)+\alpha J_{\ell}(x)v$
$\approx o(x,v,\alpha).$
Step 5: Infinitely many distinct solutions
Since $\dim(\mathcal{N})\geq 1$, the null space contains infinitely many directions. For any $v_{0}\in\mathcal{N}\setminus\{0\}$ and any $\beta\in\mathbb{R}$:
@@
generates infinitely many observationally equivalent vectors. Furthermore, if $\beta$ is chosen such that $v^{\prime}_{\beta}\not\propto v$ (which is always possible unless $v\propto v_{0}$), these vectors are geometrically distinct.
Step 6: Non-proportionality
To ensure $v^{\prime}_{\beta}\not\propto v$, we need $v+\beta v_{0}\neq cv$ for any scalar $c$. This fails only if:
$\beta v_{0}=(c-1)v$ (20)
which requires $v\in\text{span}(v_{0})$. Since $v_{0}$ is an arbitrary element of a $\dim(\mathcal{N})$-dimensional space and $v$ is fixed, this occurs with probability zero. Therefore, for generic $v$ and generic $v_{0}\in\mathcal{N}$, we have $v^{\prime}_{\beta}\not\propto v$ for almost all $\beta$.
Conclusion (Null-space mechanism)
@@
Under the local linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
B.2 Weight Space Gauge Symmetry (Supplementary Observation)
While Proposition 1 fixes model weights and asks whether v is recoverable from behavior, this section offers a complementary perspective: even relaxing the weight-fixing assumption, non-identifiability persists via linear reparameterization symmetries.
@@ B.2 Weight Space Gauge Symmetry (Supplementary Observation)
Setup: We record a supplementary observation about gauge symmetry in neural network parameterization. This is distinct from the null-space non-identifiability of Proposition 1: rather than fixing model weights and asking whether $v$ is recoverable from behavioral observations, we ask whether different $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs can produce identical input–output behavior. The answer is yes, via the linear reparameterization below. This symmetry is not realizable as a steering operation on a frozen model as it requires modifying $W_{\ell+1}$.
Step 1: Representation reparameterization
Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Define the reparameterized representation:
$h^{\prime}_{\ell}(x)=T\bigl(h_{\ell}(x)\bigr).$ (21)
If the subsequent layers can be rewritten as:
$F_{\ell\to L}(h_{\ell})=F^{\prime}_{\ell\to L}(T(h_{\ell}))$ (22)
then $(h_{\ell},v)$ and $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=DT(h_{\ell})\cdot v$ are observationally indistinguishable.
Step 2: Linear reparameterizations
Consider linear transformations $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$ is invertible. For layer $\ell+1$ with weight matrix $W_{\ell+1}\in\mathbb{R}^{d\times d}$ and bias $b_{\ell+1}$:
Original computation:
$h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ (23)
Reparameterized computation:
$h^{\prime}_{\ell+1}=\sigma(W^{\prime}_{\ell+1}h^{\prime}_{\ell}+b^{\prime}_{\ell+1})$ (24)
where $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$ and $b^{\prime}_{\ell+1}=b_{\ell+1}$.
Note: The bias remains invariant under linear reparameterization through the origin. For more general affine transformations $T(h)=Ah+c$ with translation $c\neq 0$, the bias would also transform as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$. We restrict to origin-preserving transformations for simplicity, as these suffice to establish non-identifiability. Then:
$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (25)
$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
$=h_{\ell+1}.$
Step 3: Steering under reparameterization
Original steering:
$\tilde{h}_{\ell}=h_{\ell}+\alpha v$ (26)
$\tilde{h}_{\ell+1}=\sigma(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1})$ (27)
Reparameterized steering with $v^{\prime}=Av$:
$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (28)
$=Ah_{\ell}+\alpha Av$
$=A(h_{\ell}+\alpha v)$
$=A\tilde{h}_{\ell}$.
$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$ (29)
$=\sigma\!\left(W_{\ell+1}A^{-1}A(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
$=\sigma\!\left(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
$=\tilde{h}_{\ell+1}.$
Step 4: Infinitely many reparameterizations
For any invertible matrix $A$ with $A\neq cI$ (i.e., not a scalar multiple of identity), we obtain $v^{\prime}=Av\not\propto v$. The space of such matrices has dimension $d^{2}-1$ (excluding scalar multiples), providing infinitely many distinct reparameterizations.
Conclusion (Weight-space gauge symmetry)
The symmetries of neural network parameterization imply the existence of infinitely many $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs that yield identical input–output behavior.
Appendix C Intuitive Explanations
C.1 Geometric Intuition for Null-Space Ambiguity
Visual analogy:

### 01 Apr 2026 v4

@@ 8 Limitations
Our empirical validation covers two models at mid-network layers across five semantic traits. While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions. The multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions but do not test whether explicit structural interventions such as independence constraints (Hyvärinen and Pajunen, 1999), sparsity regularization or invariance objectives (Arjovsky et al., 2019) can recover identifiable representations. We also do not directly verify that the semantic null-space estimated via activation SVD coincides with the Jacobian null-space assumed in Proposition 1; this connection is supported by the behavioral results but not formally proven.
@@ 9 Conclusion
We have shown that steering vectors are fundamentally non-identifiable under white-box single-layer access: infinitely many geometrically distinct directions produce equivalent behavioral effects due to the large null space of the activation-to-output Jacobian. This non-identifiability persists across models, traits, extraction methods and prompt distributions. The implication is that behavioral equivalence is not a sufficient basis for interpretability claims about internal representations, and that alignment methods relying on steering vectors require structural constraints beyond behavioral testing to be trusted. More broadly, any intervention-based interpretability method that validates representational claims solely through output behavior faces the same fundamental limitation.
@@ Acknowledgments
We thank Vast.ai for GPU compute resources. We acknowledge the HuggingFace platform and maintainers of the open-source models used in this work.
Ethics Statement
This work studies the theoretical and empirical foundations of activation steering in large language models. We do not introduce new steering methods or systems that could enable misuse. Our findings only highlight limitations of existing methods. We use publicly available models and standard evaluation datasets throughout and have adhered to their respective terms of use.
Disclosure of AI Usage
Large language models were used for language editing of the draft and code assistance in this work. All research ideas, technical content, experimental design, results and conclusions are the original work of the authors.
Appendix A Null-Space Dimensionality
We estimate the dimensionality of the semantically inert subspace at three layers (L/4, L/2, 3L/4) of each model. Hidden activations are collected over 200 in-distribution prompts; we compute the activation covariance matrix and apply SVD. Effective rank is defined as the number of singular values exceeding $\varepsilon=0.01$. The null-space fraction is $(d-\text{rank})/d$.
Table 4: Null-space dimensionality across three layers for both models.
Model Layer Metrics
$d$ Eff. rank NF
Llama L/4 4096 57 0.943
L/2 4096 59 0.941
3L/4 4096 63 0.937
Qwen L/4 2048 139 0.861
L/2 2048 127 0.873
3L/4 2048 116 0.884
The singular value spectra at three layers are shown in Figures 6(a) and 6(b). The surface drops sharply after the first few dimensions, with singular values collapsing to near zero beyond dimension 50–140 depending on the model. This cliff-like decay persists consistently across all three layers, indicating that the low-rank structure is not a property of any particular depth but a stable geometric feature of both architectures. The vast flat region beyond the initial peak represents the semantically inert null space: directions in this region produce no measurable change in output logits and are therefore undetectable by any behavioral evaluation. For Qwen, the decay is steeper and the plateau lower which is consistent with its smaller hidden dimension ($d=2048$) relative to Llama ($d=4096$).
Figure 6: Singular value spectra across layers. (a) Qwen and (b) Llama both exhibit a sharp cliff after the first 50–140 dimensions followed by a near-zero plateau, confirming that the vast majority of the activation space is semantically inert across all layers.
Appendix B Logit-Level Equivalence Test
To confirm that equivalence is not an artifact of semantic classifier choice, we measure raw logit distances. For each model and trait, we compute the $\ell_{2}$ distance between the next-token logit vectors produced under three conditions: (i) steering with $v$, (ii) steering with $v+v_{\perp}$ where $v_{\perp}$ is a norm-matched orthogonal direction, and (iii) steering with a norm-matched random direction $v_{\text{rand}}$. The ratio
@@ Appendix B Logit-Level Equivalence Test
quantifies how much smaller the logit shift from an orthogonal perturbation is compared to a random one. Values $R<1$ indicate that orthogonal perturbations are systematically less disruptive than random directions.
Table 5: Logit-level equivalence. Baseline = $\|\ell(0)-\ell(v)\|_{2}$; $v_{\perp}$ and Random denote $\|\ell(v)-\ell(v+v_{\perp})\|_{2}$ and $\|\ell(v)-\ell(v_{\text{rand}})\|_{2}$ respectively. $R<1$ in all conditions.
Model Trait Metrics
Baseline $v_{\perp}$ Random $R$
Qwen Formality 43.73 40.93 57.78 0.71
Sentiment 67.56 62.79 86.21 0.73
Truthfulness 113.06 118.24 165.91 0.71
Politeness 53.14 52.95 73.07 0.72
Agreeableness 75.66 64.64 96.92 0.67
Llama Formality 119.12 64.36 136.58 0.47
Sentiment 87.67 61.84 100.94 0.61
Truthfulness 206.04 121.33 226.73 0.54
Politeness 116.88 98.70 144.28 0.68
Agreeableness 119.98 85.01 128.02 0.66
$R$ ranges from $0.47$ to $0.73$ across all conditions, as reported in Table 5. Orthogonal perturbations induce logit shifts that are 27–53% smaller than random directions of the same norm. In several Llama conditions, the orthogonal perturbation logit distance is also smaller than the baseline distance, indicating that $v_{\perp}$ moves the model less than $v$ itself does. These results confirm that the equivalence observed in semantic scores is not a property of the classifiers: it is present directly in the model’s output distribution. The full bar-chart comparisons are shown in Figures 7(a) and 7(b).
Figure 7: Logit $\ell_{2}$ distances for (a) Qwen and (b) Llama. Each group shows three bars per trait: Baseline ($\|\ell(0)-\ell(v)\|_{2}$), $v_{\perp}$ ($\|\ell(v)-\ell(v+v_{\perp})\|_{2}$) and Random ($\|\ell(v)-\ell(v_{\text{rand}})\|_{2}$). The $v_{\perp}$ bar is consistently shorter than Random across all traits which confirms that orthogonal perturbations are less disruptive to the output distribution than unstructured directions of equal norm.
Appendix C Detailed Proof of Proposition 1
We provide a constructive proof of Proposition 1, formalizing the null-space ambiguity argument presented in Section 5.
Setup.
Consider the local linear approximation of the steering effect at layer $\ell$:
@@ Appendix C Detailed Proof of Proposition 1
where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian of the output logits with respect to the layer-$\ell$ activations, $V$ is the vocabulary size, and $d$ is the hidden dimension.
C.1 Null-space characterization
@@
C.2 Rank bound and effective rank
The Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ satisfies $\operatorname{rank}(J_{\ell})\leq\min(V,d)$. In modern language models $V\approx 50{,}000$ and $d\approx 2{,}000$8,0008{,}000, so the theoretical maximum rank is $d$. However, empirical evidence indicates that output distributions lie on a low-dimensional manifold. The effective rank, defined as
$\operatorname{rank}_{\varepsilon}(J_{\ell})=\#\{\sigma_{i}:\sigma_{i}>\varepsilon\cdot\sigma_{\max}\}$ (14)
where $\sigma_{i}$ are singular values of $J_{\ell}$ and $\varepsilon$ is a threshold (e.g., $10^{-2}$), is substantially lower than $d$ in practice. Our measurements (Appendix A) yield effective ranks of 57 and 139 for Llama and Qwen respectively, leaving null-space fractions of 94% and 86%. The constructive argument below requires only $\dim(\mathcal{N})\geq 1$.
C.3 Constructing observationally equivalent vectors
For any steering vector $v\in\mathbb{R}^{d}$ and any $v_{0}\in\mathcal{N}$, define $v^{\prime}=v+v_{0}$. Then for all prompts $x$ and all steering strengths $\alpha$:
$J_{\ell}(x)v^{\prime}=J_{\ell}(x)(v+v_{0})=J_{\ell}(x)v+J_{\ell}(x)v_{0}=J_{\ell}(x)v$ (15)
where the last step uses $v_{0}\in\ker(J_{\ell})$. Therefore:
$o(x,v^{\prime},\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}=o(x,0,0)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha)$ (16)
C.4 Infinitely many distinct solutions
Since $\dim(\mathcal{N})\geq 1$, the null space contains infinitely many directions. For any $v_{0}\in\mathcal{N}\setminus\{0\}$ and any $\beta\in\mathbb{R}$, the family
@@ C.4 Infinitely many distinct solutions
generates infinitely many observationally equivalent vectors. To ensure $v^{\prime}_{\beta}\not\propto v$, we need $v+\beta v_{0}\neq cv$ for any scalar $c$, i.e., $\beta v_{0}\neq(c-1)v$, which fails only when $v\in\operatorname{span}(v_{0})$. For generic $v$ and $v_{0}\in\mathcal{N}$, this condition holds with probability zero. Therefore, for almost all $\beta\neq 0$, the vector $v^{\prime}_{\beta}$ is geometrically distinct from $v$ yet observationally equivalent.
Conclusion.
@@
Under the local linear approximation and whenever $\operatorname{rank}(J_{\ell})<d$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent to $v$. Proposition 1 follows. $\square$
C.5 Scope of the linearization
The proof relies on the approximation in Equation 4. This holds exactly for piecewise-linear activations (e.g., ReLU MLPs, on activation-region interiors) and as a first-order approximation for smooth nonlinearities (GeLU, attention softmax) with error $O(\alpha^{2})$. For the steering strengths used in our experiments ($\alpha\leq 3$), the nonlinear residual is empirically negligible: the scale-invariance results in Section 6.4 confirm that equivalence holds across the full operationally relevant range without degradation. Extending the proof to the exact nonlinear regime would require characterizing the Hessian structure of $F_{\ell\to L}$ and is left for future work.
Appendix D Weight-Space Gauge Symmetry
@@
This section records a supplementary observation about gauge symmetry in neural network parameterization. It is distinct from the null-space argument of Proposition 1: rather than fixing model weights and asking whether $v$ is recoverable from behavioral observations, we ask whether different weight-vector pairs $(W,v)$ and $(W^{\prime},v^{\prime})$ can produce identical input–output behavior. The answer is yes, via the linear reparameterization argument below. This symmetry is not realizable as a steering operation on a frozen model since it requires modifying $W_{\ell+1}$. We include it because it establishes a complementary non-identifiability result at the level of model parameterization.
Representation reparameterization.
Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and define the reparameterized representation $h^{\prime}_{\ell}(x)=T(h_{\ell}(x))$. If the subsequent layers satisfy $F_{\ell\to L}(h_{\ell})=F^{\prime}_{\ell\to L}(T(h_{\ell}))$, then the pair $(h_{\ell},v)$ and $(h^{\prime}_{\ell},v^{\prime})$ with $v^{\prime}=DT(h_{\ell})\cdot v$ are observationally indistinguishable. We now show this holds for linear $T$.
Linear reparameterization.
Restrict to $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$ is invertible, so $h^{\prime}_{\ell}(x)=Ah_{\ell}(x)$. For layer $\ell+1$ with weight matrix $W_{\ell+1}$ and bias $b_{\ell+1}$, define the reparameterized weights $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$ and $b^{\prime}_{\ell+1}=b_{\ell+1}$. The bias is invariant because we restrict to origin-preserving transformations; for affine $T(h)=Ah+c$ the bias transforms as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$, but linear $T$ suffices to establish non-identifiability. Under this reparameterization:
$h^{\prime}_{\ell+1}=\sigma(W^{\prime}_{\ell+1}h^{\prime}_{\ell}+b^{\prime}_{\ell+1})=\sigma(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1})=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})=h_{\ell+1}$ (18)
so the reparameterized network produces identical activations at layer $\ell+1$ and all subsequent layers.
Steering under reparameterization.
Under the original parameterization, steering applies $\tilde{h}_{\ell}=h_{\ell}+\alpha v$, giving $\tilde{h}_{\ell+1}=\sigma(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1})$. Under the reparameterization with $v^{\prime}=Av$, the steered representation at layer $\ell$ is:
$\tilde{h}^{\prime}_{\ell}=h^{\prime}_{\ell}+\alpha v^{\prime}=Ah_{\ell}+\alpha Av=A(h_{\ell}+\alpha v)=A\tilde{h}_{\ell}$ (19)
and the next-layer activation is:
$\tilde{h}^{\prime}_{\ell+1}=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)=\sigma\!\left(W_{\ell+1}A^{-1}A(h_{\ell}+\alpha v)+b_{\ell+1}\right)=\tilde{h}_{\ell+1}$ (20)
so the steered activation at layer $\ell+1$ is identical under both parameterizations.
Infinitely many reparameterizations.
For any invertible $A$ with $A\neq cI$, the steering vectors $v^{\prime}=Av$ and $v$ are geometrically distinct ($v^{\prime}\not\propto v$) yet produce identical outputs. The space of invertible matrices with $A\neq cI$ has dimension $d^{2}-1$, providing infinitely many such reparameterizations. Non-identifiability therefore persists even when model weights are not fixed: infinitely many $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs yield identical input–output behavior, confirming that the ambiguity is a structural property of the parameterization rather than a limitation of any particular extraction method.
Appendix E Mathematical Derivations
E.1 Dimension of the Observational Equivalence Class
For a steering vector $v\in\mathbb{R}^{d}$ and Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with rank $r$, two vectors $v$ and $v^{\prime}$ are observationally equivalent if and only if $J_{\ell}v=J_{\ell}v^{\prime}$, i.e., $J_{\ell}(v^{\prime}-v)=0$, i.e., $v^{\prime}-v\in\ker(J_{\ell})$. The equivalence class is therefore the affine subspace:
$v]=\{v^{\prime}\in\mathbb{R}^{d}:J_{\ell}v^{\prime}=J_{\ell}v\}=v+\ker(J_{\ell})$ (21)
By the rank-nullity theorem, $\dim(\ker(J_{\ell}))=d-r$. Let $\{u_{1},\ldots,u_{d-r}\}$ be an orthonormal basis for $\ker(J_{\ell})$ obtained via SVD of $J_{\ell}$. Then:
$v]=\left\{v+\sum_{i=1}^{d-r}\beta_{i}u_{i}:\beta_{i}\in\mathbb{R}\right\}$ (22)
The measure of this set in $\mathbb{R}^{d}$ is:
$\text{vol}([v])=\int_{\mathbb{R}^{d-r}}d\beta_{1}\cdots d\beta_{d-r}=\infty$ (23)
so the equivalence class has infinite volume in $\mathbb{R}^{d}$ whenever $r<d$. Additionally, $v$ and $cv$ for $c\neq 0$ produce equivalent outputs up to rescaling $\alpha\to c\alpha$, so the equivalence class modulo scaling is a (dr)(d-r)-dimensional projective space $\mathbb{P}^{d-r}$. The practical implication is that steering vector recovery is a severely underdetermined inverse problem: there are infinitely many solutions consistent with any finite set of behavioral observations, parameterized by the free coefficients $\beta_{1},\ldots,\beta_{d-r}$.
To quantify the degree of under-determination, define the non-identifiability ratio:
$\rho=\frac{d-r}{r}=\frac{\dim(\ker(J_{\ell}))}{\operatorname{rank}(J_{\ell})}$ (24)
which measures unidentifiable directions per identifiable one. For Llama-3.1-8B with $d=4096$ and $r=57$:
$\rho_{\text{Llama}}=\frac{4096-57}{57}=\frac{4039}{57}\approx 70.9$ (25)
For Qwen2.5-3B with $d=2048$ and $r=139$:
$\rho_{\text{Qwen}}=\frac{2048-139}{139}=\frac{1909}{139}\approx 13.7$ (26)
In both cases $\rho\gg 1$: the unobservable dimensions vastly outnumber the observable ones. The null space is not a thin subspace but the dominant structure of the ambient geometry.
E.2 Fisher Information and Cramér-Rao Bound
We formalize why collecting more behavioral data cannot resolve the null-space ambiguity. Consider the linear Gaussian observation model for a single prompt $x$:
$o(x)=J_{\ell}(x)v+\eta,\quad\eta\sim\mathcal{N}(0,\sigma^{2}I_{V})$ (27)
where $o(x)\in\mathbb{R}^{V}$ are the observed output logits and $v\in\mathbb{R}^{d}$ is the parameter to be estimated. The log-likelihood is:
$\log p(o\mid x,v)=-\frac{1}{2\sigma^{2}}\|o-J_{\ell}(x)v\|^{2}+\text{const}$ (28)
The score function is $\frac{\partial\log p}{\partial v}=\frac{1}{\sigma^{2}}J_{\ell}(x)^{\top}(o-J_{\ell}(x)v)$, and the Fisher information matrix is:
$\mathcal{I}(v)=\mathbb{E}_{o}\left[\frac{\partial\log p}{\partial v}\frac{\partial\log p}{\partial v}^{\top}\right]=\frac{1}{\sigma^{2}}J_{\ell}(x)^{\top}J_{\ell}(x)$ (29)
The Cramér-Rao lower bound states that for any unbiased estimator $\hat{v}$:
$\text{Cov}(\hat{v})\succeq\mathcal{I}(v)^{-1}=\sigma^{2}\bigl(J_{\ell}(x)^{\top}J_{\ell}(x)\bigr)^{+}$ (30)
where $(\cdot)^{+}$ is the Moore-Penrose pseudoinverse. For a null-space direction $u_{i}\in\ker(J_{\ell})$, we have $J_{\ell}u_{i}=0$, so:
$u_{i}^{\top}\mathcal{I}(v)u_{i}=\frac{1}{\sigma^{2}}u_{i}^{\top}J_{\ell}^{\top}J_{\ell}u_{i}=\frac{1}{\sigma^{2}}\|J_{\ell}u_{i}\|^{2}=0$ (31)
Since $\mathcal{I}(v)$ is singular in the direction $u_{i}$, its pseudoinverse satisfies $u_{i}^{\top}\mathcal{I}(v)^{+}u_{i}=\infty$ (the pseudoinverse assigns zero eigenvalue to $u_{i}$, meaning the bound is vacuously infinite). Therefore:
$\text{Var}(u_{i}^{\top}\hat{v})\geq u_{i}^{\top}\mathcal{I}(v)^{+}u_{i}=\infty\quad\forall\,u_{i}\in\ker(J_{\ell})$ (32)
Multi-prompt extension.
Stacking observations across $N$ prompts $\{x_{1},\ldots,x_{N}\}$ gives:
$\mathbf{o}=J_{\text{stack}}v+\boldsymbol{\eta},\quad J_{\text{stack}}=\begin{bmatrix}J_{\ell}(x_{1})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}\in\mathbb{R}^{(NV)\times d}$ (33)
with aggregate Fisher information $\mathcal{I}_{N}(v)=\frac{1}{\sigma^{2}}J_{\text{stack}}^{\top}J_{\text{stack}}=\frac{1}{\sigma^{2}}\sum_{i=1}^{N}J_{\ell}(x_{i})^{\top}J_{\ell}(x_{i})$. The null space of $\mathcal{I}_{N}(v)$ is:
$\ker(\mathcal{I}_{N})=\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$ (34)
Any direction $u\in\bigcap_{i}\ker(J_{\ell}(x_{i}))$ satisfies $J_{\ell}(x_{i})u=0$ for all $i$, meaning it contributes zero information regardless of $N$. If the intersection of null spaces is non-trivial — which empirically holds when all prompts probe the same effective output subspace — then $\ker(\mathcal{I}_{N})\neq\{0\}$ for all $N$ and the Cramér-Rao bound remains infinite along those directions. Non-identifiability is therefore not a small-sample problem: it is an intrinsic property of the model geometry that no amount of behavioral data can resolve.
Appendix F Intuitive Explanations
F.1 Geometric intuition for null-space ambiguity